The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

                                    mcnews@astro.uiuc.edu
                         http://www.astro.uiuc.edu/mcnews/MCNews.html
                            http://www.astro.uni-bonn.de/~mcnews/

No. 29 February 3, 1999

Contents

Abstracts of 10 refereed papers  1
Abstract of a non-refereed paper 8
Thesis abstract  9
Job opportunities  10
Meeting announcements  11

Abstracts of Refereed Papers

Molecular abundance variations in the Magellanic Clouds

A. Heikkilä^{1,3}, L.E.B. Johansson^{1}, and H. Olofsson^{2}

1Onsala Space Observatory, S–439 92 Onsala, Sweden 2Stockholm Observatory, S–133 36 Saltsjöbaden, Sweden 3present address: Observatory, P.O. Box 14, FIN–000 14 University of Helsinki, Finland

We have observationally studied the effect of metallicity and far-ultraviolet (FUV) radiation on the physical conditions and the molecular abundances in interstellar clouds in the Small and the Large Magellanic Clouds (SMC and LMC, respectively). Spectral line emission from a number of molecules was observed in a sequence of clouds with positions in and between the 30 Doradus (30Dor) and the southern part of the N159 region in the LMC, and in one cloud (N27, also denoted LIRS 49) located in the SMC bar. Physical conditions and molecular abundances were estimated from the observational data by excitation and radiative transfer calculations. A comparison of the molecular abundances in clouds in the SMC, the LMC, and the Galaxy is presented. We also report the first detection of hydrogen sulphide (ortho-H_{2}S) in an extragalactic source, detections of methanol (CH_{3}OH) in thermal emission and methyl acetylene (CH_{3}CCH), and a tentative detection of thio-formaldehyde (H_{2}CS) in N159W.

The abundances (relative to H_{2}) of molecular species (except CO) in the LMC sources and in N27 are estimated to be typically 5×10^{−10}, and 1×10^{−10}, respectively. These values apply to the gas volume defined by the CO line emission. Relative to Galactic clouds, the abundances in N159W (our reference cloud) are five to twenty times lower. In two of the clouds: N27 and the centremost cloud in 30Dor (30Dor-10), the derived abundances deviate significantly from those in the other clouds in our sample, by being on the average six and eight times lower, respectively. In N27, the most likely explanation is the lower metallicity in the SMC, whereas the underabundance in 30Dor-10 is probably mainly caused by a more rapid photodissociation due to the more intense FUV radiation in this area. An alternative explanation for the underabundances in both N27 and 30Dor-10 would be a higher H/H_{2} ratio inside these molecular clouds. The ethynyl radical (C_{2}H), with an estimated average abundance of 5×10^{−9} in seven clouds in the LMC and 3×10^{−9} in two clouds in the SMC, is the most abundant observed trace molecule after the CO isotopomers. Qualitatively, the high C_{2}H abundance can be explained as reflecting the C^{+}-rich and FUV photon-rich environment, i.e., a chemistry characteristic for photon-dominated regions.

For N27 we have, using HCO^{+} and H^{13}CO^{+} data, estimated the gas-phase ^{12}C/^{13}C ratio to be 40–90, a range that encompasses the values found in N159W and in Galactic disc clouds.

In all clouds in our sample, the number density estimates from an excitation analysis of CS, SO, HCO^{+}, HCN and H_{2}CO are in the range (1–100)×10^{4} cm^{−3}. CO data gives a lower limit of a few×10^{3} cm^{−3}. However, the average densities (estimated from the virial mass) are significantly lower, typically a few×10^{2} cm^{−3}, suggesting that the clouds (as probed by trace molecules) are very clumpy with volume-filling factors ≪1. In N159W, where our data-base is by far most extensive, the number density and the kinetic temperature of molecular hydrogen in the dense part of the gas are estimated to be (1–10)×10^{5} cm^{−3} and 25±10 K, respectively. The corresponding numbers in N27 are, although based on less data than in N159W, (5–50)×10^{4} cm^{−3} and 15±5 K. Thus, the metallicity difference between the LMC and the SMC does not seem to affect the density and the temperature of the gas dramatically.

In the SMC, the CO(J=1–0)/HCO^{+}(J=1–0) line intensity ratio follows the same trend with respect to the star-formation activity as in the LMC: a lower ratio is found in clouds with a more vigorous star-formation activity. A similar trend is also exhibited by the CO(J=1–0)/C_{2}H(N=1–0) line intensity ratio in the LMC.

Accepted by: Astronomy & Astrophysics, Main Journal

For preprints, contact arto@obs13.astro.helsinki.fi

Bok Globules in the LMC

D.R. Garnett^{1,2}, J.R. Walsh^{3}, Y.-H. Chu^{4}, and B.M. Lasker^{5}

1 Astronomy Department, University of Minnesota, USA 2 Current address: Steward Observatory, University of Arizona, USA 3 Space Telescope European Coordinating Facility, European Southern Observatory, Germany 4 Department of Astronomy, University of Illinois, USA 5 Space Telescope Science Institute, USA

We report the discovery of small, isolated dust clouds in the Large Magellanic Cloud, which are excellent candidates for counterparts to the Bok globules observed in the Galaxy. We detect these small clouds silhouetted against diffuse Hα emission, based on parallel imaging with the WFPC-2 on HST. The clouds we identify as Bok globule candidates have typical sizes of approximately 1′′, corresponding to about 0.25 pc linear diameter at the distance of the LMC. We derive lower limits to the optical depth within the dark clouds, and masses assuming that the clouds have density distributions similar to Galactic Bok globules. The sizes and estimated masses for LMC globules are comparable to those estimated for Galactic globules. An extended sample of such objects would be excellent targets for high-resolution infrared and millimeter observations to study low-mass star formation in such clouds in low-metallicity environments, where the distance is well known.

Accepted by: The Astronomical Journal

For preprints, contact dgarnett@as.arizona.edu

Very young massive stars in the Small Magellanic Cloud, revealed by HST

M. Heydari-Malayeri^{1}, M.R. Rosa^{2}, H. Zinnecker^{3}, L. Deharveng^{4}, and V. Charmandaris^{1}

1 demirm, Observatoire de Paris, 61 Avenue de l’Observatoire, F-75014 Paris, France 2 Space Telescope European Coordinating Facility, ESO, Karl-Schwarzschild-Str. 2, D-85748 Garching, Germany 3 Astrophysikalisches Institut Potsdam, An der Sternwarte 16, D-14482 Potsdam, Germany 4 Observatoire de Marseille, 2 Place Le Verrier, F-13248 Marseille Cedex 4, France

High spatial resolution imaging with the Hubble Space Telescope allowed us to resolve the compact H ii region N81 lying in the Small Magellanic Cloud (SMC). We show the presence of a tight cluster of newborn massive stars embedded in this nebular “blob” of ∼ 10′′ across. This is the first time the stellar content and internal morphology of such an object is uncovered. These are among the youngest massive stars in this galaxy accessible to direct observations at ultraviolet and optical wavelengths. Six of them are grouped in the core region of ∼ 2′′ diameter, with a pair of the main exciting stars in the very center separated by only 0′′.27 or 0.08 pc. The images display violent phenomena such as stellar winds, shocks, ionization fronts, typical of turbulent starburst regions. Since the SMC is the most metal-poor galaxy observable with very high angular resolution, these observations provide important templates for studying star formation in the very distant metal-poor galaxies which populate the early Universe.

Accepted by: Astronomy and Astrophysics For preprints, contact heydari@obspm.fr

Also available from the URL http://wwwusr.obspm.fr/^{∼}heydari/projects/N81

Supergiant Abundances in the SMC: Probes of Evolution

K.A. Venn^{1}

1 Macalester College, Dept of Physics & Astronomy, 1600 Grand Ave., St. Paul, MN, 55105, USA

New abundances of N, O, Na, Mg, Si, Ca, Sc, Ti, Cr, Fe, Sr, Zr, and Ba are presented for 10 Atype supergiants in the SMC, plus upper limits for C. In interpreting the CNO results for constraints on stellar evolution theories, careful attention has been paid to the comparison abundances, i.e., the present day abundances of SMC nebulae and B-dwarf stars. These new results are also compared to published results from F-K supergiant analyses, and found to be in good agreement when both sets of data are carefully examined as differential (SMC minus Galactic standard) abundances.

With the exception of nitrogen, very small star-to-star abundance variations are found for all elements in this analysis. The N variations are not predicted by standard stellar evolution models. Instead, the results support the new predictions reported from rotating stellar models, where the range in nitrogen is the result of partial mixing of CN-cycled gas from the stellar interior due to main-sequence rotation at different rates (c.f., Langer & Heger 1998). The overall overabundance of nitrogen in the sampled stars also implies these stars have undergone the first dredge-up in addition to having been mixed while on the main-sequence.

The alpha-elements (O, Mg, Si, Ca, Ti) have similar underabundances to Fe, which is not the same as seen in metal-poor stars in the solar neighborhood of the Galaxy. In addition, certain light s-process elements (Zr, Ba) are slightly more underabundant than Fe, which is predicted by the bursting chemical evolution model presented by Pagel & Tautvais̆ienė (1998) for the SMC.

Accepted by: The Astrophysical Journal

For preprints, contact venn@clare.physics.macalester.edu
Also available from the URL http://xxx.lanl.gov/ps/astro-ph/9901306

Period - luminosity - color - radius relationships of Cepheids as a function of metallicity: evolutionary effects.

Yann Alibert^{1}, Isabelle Baraffe^{1,2}, P. Hauschildt^{3} and France Allard^{1}

1 Ecole Normale supérieure-CRAL, Lyon, France 2 Max-Planck Institut für Astrophysik, Garching, Germany 3 Dept. of Physics and Astronomy and Center for Simulational Physics, University of Georgia Athens, USA

Based on consistent evolutionary and pulsation calculations, we analyse the effect of metallicity and of different convection treatments in the stellar models on period - magnitude, - color and - radius relationships. In order to perform an accurate comparison with observations, we have computed grids of atmosphere models and synthetic spectra for different metallicities, covering the range of effective temperatures and gravities relevant for Cepheids. The models are compared to recent observations of galactic and Magellanic Clouds Cepheids. Unprecedented level of agreement is found between models and observations.

We show that within the range of metallicity for the Galaxy and the Magellanic Clouds, a change of slope in the period - luminosity (PL) relationship is predicted at low periods, due to the reduction of the blue loop during core He burning. The minimum mass undergoing a blue loop and consequently the critical period at which this change of slope occurs depend on the metallicity Z and on the convection treatment in the stellar models.

However, besides this change of slope, we do not find any significant effect of metallicity on period - magnitude relationships from V to K bands, and on period - color relationships in IR colors. We only find a detectable effect of Z on (B − V) colors. These results are not affected by uncertainties inherent to current stellar models, mainly due to convection treatment.

Accepted by: Astronomy and Astrophysics

For preprints, contact ibaraffe@ens-lyon.fr
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9901294
or by anonymous ftp at ftp.ens-lyon.fr/pub/users/CRAL/ibaraffe/cepheid/ABHA98.ps

On the nature of the peculiar hot star in the young LMC cluster NGC 1818

James Liebert^{1}

1 Steward Observatory, Univ. of Arizona and Space Telescope Science Institute, USA

The blue star reported in the field of the young LMC cluster NGC 1818 by Elson et al. (1998) has the wrong luminosity and radius to be a “luminous white dwarf” member of the cluster. In addition, unless the effective temperature range quoted by the authors is a drastic underestimate, the luminosity is much too low for it to be a cluster member in the post-AGB phase. Other possibilities, including that of binary evolution, are briefly discussed. However, the implication that the massive main sequence turnoff stars in this cluster can produce white dwarfs (instead of neutron stars) from single-star evolution needs to be reconsidered.

Accepted by: Astrophysical Journal Letters For preprints, contact liebert@stsci.edu

Also available from http://xxx.lanl.gov/abs/astro-ph/9901312

A secondary clump of red giant stars: why and where

Léo Girardi

Max-Planck-Institut für Astrophysik, Karl-Schwarzschild-Str. 1, D-85740 Garching bei München, Germany

Based on the results of detailed population synthesis models, Girardi et al. (1998) recently claimed that the clump of red giants in the colour–magnitude diagram (CMD) of composite stellar populations should present an extension to lower luminosities, which goes down to about 0.4 mag below the main clump. This feature is made of stars just massive enough for having ignited helium in non-degenerate conditions, and therefore corresponds to a limited interval of stellar masses and ages. In the present models, which include moderate convective overshooting, it corresponds to ∼ 1 Gyr old populations.

In this paper, we go into more details about the origin and properties of this feature. We first compare the clump theoretical models with data for clusters of different ages and metallicities, basically confirming the predicted behaviours. We then refine the previous models in order to show that:

(i) The faint extension is expected to be clearly separated from the main clump in the CMD of metalrich populations, defining a ‘secondary clump’ by itself. (ii) It should be present in all galactic fields containing ∼ 1 Gyr old stars and with mean metallicities higher than about Z = 0.004. (iii) It should be particularly strong, if compared to the main red clump, in galaxies which have increased their star formation rate in the last Gyr or so of their evolution. In fact, secondary clumps similar to the model predictions are observed in the CMD of nearby stars from Hipparcos data, and in those of some Large Magellanic Cloud fields observed to date. There are also several reasons why this secondary clump may be missing or hidden in other observed CMDs of galaxy fields. For instance, it becomes indistinguishable from the main clump if the photometric errors or differential absorption are larger than about 0.2 mag. Nonetheless, this structure may provide important constraints to the star formation history of Local Group galaxies. We comment also on the intrinsic luminosity variation and dispersion of clump stars, which may limit their use as either absolute or relative distance indicators, respectively.

Submitted to: Monthly Notices of the Royal Astronomical Society

For preprints, contact leo@mpa-garching.mpg.de

Also available from the URL astro-ph/9901319

Zeroing the Stellar Isochrone Scale: The Red Giant Branch Clump Luminosity at Intermediate Metallicity

B.A. Twarog^{1}, B.J. Anthony-Twarog^{1} and A.R. Bricker^{1}

1 Univ. of Kansas, Dept. of Physics and Astronomy, Lawrence, KS 66045-2151, USA

The color-magnitude diagrams of the open clusters NGC 2420 and NGC 2506 have been investigated as intermediate links between the solar neighborhood and the Magellanic Clouds. Two sets of theoretical isochrones which include convective overshoot are zeroed to the sun at solar abundance and to the unevolved main sequence dwarfs of the Hipparcos catalog at [Fe/H] = –0.4, requiring a differential of 0.4 mag between the unevolved main sequences at a given color. Adopting E(B − V) = 0.04 and [Fe/H] = –0.39 for NGC 2506 and E(B − V) = 0.04 and [Fe/H] = –0.29 for NGC 2420, the respective apparent moduli are (m − M) = 12.70 and 12.15, while the ages of both clusters are approximately 1.9 ± 0.2 Gyr or 2.2 ± 0.2 Gyr, depending on the choice of isochrones. From the composite giant branch of the two clusters, the mean clump magnitudes in V and I are found to be +0.47 and –0.48 (–0.17,+0.14), respectively. Applying a metallicity correction to the M_{I} values, the cluster sample of Udalski (1998) leads to (m − M)_{0} = 18.42 (+0.17,–0.15) and 18.91 (+0.18,–0.16) for the LMC and SMC, respectively. A caveat to this discussion and potentially to the claim that clusters of the same abundance and age are identical is the observation that the (V −I) colors of the red giants in NGC 2506 are significantly redder at a given (B − V) than the giants in clusters of comparable age and/or metallicity. The distance scale above has been derived using the general cluster relation between (B − V) and (V − I). If the CCD photometry in NGC 2506 is correctly tied to the standard system, M_{I} for the clump will decrease and the distance moduli should increase by 0.1 mag.

Accepted by: The Astronomical Journal, April 1999

Also available from the URL http://kubarb.phsx.ukans.edu/∼anthony/anthony.html
or by anonymous ftp at kubarb.phsx.ukans.edu in /pub directory

The third dredge-up and the carbon star luminosity functions in the Magellanic Clouds

Paola Marigo^{1,2}, Léo Girardi^{1} and Alessandro Bressan^{3}

1 Max-Planck-Institut für Astrophysik, Karl-Schwarzschild-Str. 1, D-85740 Garching bei München, Germany 2 Department of Astronomy, University of Padova, Vicolo dell’Osservatorio 5, I-35122 Padova, Italy 3 Astronomical Observatory, Vicolo dell’Osservatorio 5, I-35122, Padova, Italy

We investigate the formation of carbon stars as a function of the stellar mass and parent metallicity. Theoretical modelling is based on an improved scheme for treating the third dredge-up in synthetic calculations of thermally pulsing asymptotic giant branch (TP-AGB) stars. In this approach, the usual criterion (based on a constant minimum core mass for the occurrence of dredge-up, M_{c}^{min}) is replaced by one on the minimum temperature at the base of the convective envelope, T_{b}^{dred}, at the stage of the post-flash luminosity maximum. Envelope integrations then allow determination of M_{c}^{min} as a function of stellar mass, metallicity, and pulse strength (see Wood 1981), thus inferring if and when dredge-up first occurs. Moreover, the final possible shut down of the process is predicted.

Extensive grids of TP-AGB models were computed by Marigo (1998ab) using this scheme. In this paper, we present and discuss the calibration of the two dredge-up parameters (i.e. efficiency λ and T_{b}^{dred}) aimed at reproducing the carbon star luminosity function (CSLF) in the LMC, using TP-AGB models with original metallicity Z = 0.008. In addition to this, the effects of different input quantities on the theoretical CSLF are analysed. It turns out that the faint tail is almost insensitive to the history of star formation rate (SFR) in the parent galaxy, in contrast to the bright wing which may be more affected by the details of the recent history. Actually, we find that the faint end of the CSLF is essentially determined by the temperature parameter T_{b}^{dred}. Once the faint end is reproduced, the peak location is a stringent calibrator of the efficiency parameter λ. The best fit to the observed CSLF in the LMC is obtained with λ = 0.50, log T_{b}^{dred} = 6.4, and a constant SFR up to an age of about 5 × 10^{8} yr. This recent drop of the SFR is invoked to remove a slight excess of bright carbon stars otherwise predicted.

A good fit to the observed CSLF in the SMC is then easily derived from the Z = 0.004 models, with a single choice of parameters λ = 0.65, log T_{b}^{dred} = 6.4, and a constant SFR over the entire significant age interval. The result for λ is consistent with the theoretical expectation that the third dredge-up is more efficient at lower metallicities.

Accepted by: Astronomy and Astrophysics For preprints, contact marigo@pd.astro.it

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9901235

The Collisions of HVCs With a Magnetized Gaseous Galactic Disk

Alfredo Santillán^{1,2}, José Franco^{2}, Marco Martos^{2} and Jongsoo Kim^{3}

1 Cómputo Aplicado–DGSCA, Universidad Nacional Autónoma de México, A.P. 70-264, 04510 México D.F., México 2 Instituto de Astronomía, Universidad Nacional Autónoma de México, A. P. 70-264, 04510 México D.F., México 3 Korea Astronomy Observatory, San 36-1, Hwaam-Dong, Yusong-Ku, Taejon 305-348, Korea

We present two-dimensional MHD numerical simulations for the interaction of high-velocity clouds with both magnetic and non-magnetic Galactic thick gaseous disks. For the magnetic models, the initial magnetic field is oriented parallel to the disk, and we consider two different field topologies (with and without tension effects): parallel and perpendicular to the plane of motion of the clouds. The impinging clouds move in oblique trajectories and fall toward the central disk with different initial velocities. The B-field lines are distorted and compressed during the collision, increasing the field pressure and tension. This prevents the cloud material from penetrating into the disk, and can even transform a high-velocity inflow into an outflow, moving away from the disk. The perturbation creates a complex, turbulent, pattern of MHD waves that are able to traverse the disk of the Galaxy, and induce oscillations on both sides of the plane. Thus, the magnetic field efficiently transmits the perturbation over a large volume, but also acts like a shield that inhibits the mass exchange between the halo and the disk. For the non-magnetized cases, we also uncover some novel features: the evolution of the shocked layer generates a tail that oscillates, creating vorticity and turbulent flows along its trajectory.

Accepted by: The Astrophysical Journal For preprints, contact Alfredo Santillan, alfredo@astroscu.unam.mx

Abstracts of Non-Refereed Papers

Cepheid Variables in the LMC and SMC

D.L. Welch^{1}, C. Alcock^{2,3}, R.A. Allsman^{4}, D.R. Alves^{2,5,6}, T.S. Axelrod^{7}, A.C. Becker^{8}, D.P. Bennett^{2,3,9}, K.H. Cook^{2,3}, K.C. Freeman^{7}, K. Griest^{3,10}, D.W. Kurtz^{12}, M.J. Lehner^{3,10,11}, S.L. Marshall^{2,3}, D. Minniti^{2,3}, B.A. Peterson^{7}, M.R. Pratt^{3,8,13}, P.J. Quinn^{14}, A.W. Rodgers^{5,15}, A. Rorabeck^{1}, W. Sutherland^{16}, A. Tomaney^{8}, T. Vandehei^{3,10}; (The MACHO Collaboration)

1 Dept. of Physics & Astronomy, McMaster University, Hamilton, Ontario, L8S 4M1 Canada 2 Lawrence Livermore National Laboratory, Livermore, CA 94550, USA 3 Center for Particle Astrophysics, University of California, Berkeley, CA 94720, USA 4 Supercomputing Facility, Australian National University, Canberra, ACT 0200, Australia 5 Department of Physics, University of California, Davis, CA 95616, USA 6 Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 7 Mt. Stromlo and Siding Spring Observatories, Australian National University, Weston Creek, ACT 2611, Australia 8 Departments of Astronomy and Physics, University of Washington, Seattle, WA 98195, USA 9 Physics Department, University of Notre Dame, Notre Dame, IN 46556, USA 10 Department of Physics, University of California, San Diego, La Jolla, CA 92093, USA 11 Department of Physics, University of Sheffield, Sheffield, S3 7RH, U.K. 12 Department of Astronomy, University of Cape Town, Rondebosch 7701, South Africa 13 LIGO Project, MIT, Room 20B-145, Cambridge, MA 02139, USA 14 European Southern Observatory, Karl-Schwarzschild Str. 2, D-85748, Garching, Germany 15 Deceased. 16 Department of Physics, University of Oxford, Oxford OX1 3RH, U.K.

In this paper, we will review major new results regarding classical Cepheids, in the Large Magellanic Cloud (LMC) and Small Magellanic Cloud (SMC). Specifically, we discuss recent work regarding multimode Cepheids and describe new observations of a W Vir star (HV 5756) and a Cepheid which are each in eclipsing binary systems. An additional interesting pulsating supergiant in an eclipsing system is also identified. Ephemerides for eclipses for the three systems are provided.

To appear in: IAU Symposium 190, “New Views of the Magellanic Clouds”, eds. Chu et al., ASP Conf. Ser.

For preprints, contact welch@physics.mcmaster.ca
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9901282

Thesis Abstract

Young Stellar Populations in the Large Magellanic Cloud: Surprises from HST Observations

Martino Romaniello

Scuola Normale Superiore di Pisa, Italy and Space Telescope Science Institute, Baltimore, USA. Current address: European Southern Observatory, Garching bei München, Germany.

The general purpose of this Thesis is to determine in detail some important properties of young stellar populations. This is achieved by devising two new methods to analyze the data. The first is a method that allows one to recover the intrinsic temperature and luminosity, as well as the reddening, for individual stars by fitting the measured broad band magnitudes to the ones calculated with model atmospheres. The second is a procedure to infer directly (i.e. without imposing any model) the Star Formation Rate as a function of time and the underlying Initial Mass Function from the HR diagrams constructed for the observed fields. We have applied these two methods to measure ages, mass functions, and spatial distributions of stellar populations in the neighborhood of Supernova 1987A in the Large Magellanic Cloud and in a “parallel” field 120 pc away. The observations were taken with WFPC2 on board HST in 6 wide bands filters and a narrow one centered on the Hα line.

Our main conclusions are:

Spatial distribution: high (M > 6 M_{⊙}) and low (M < 2 M_{⊙}) mass stars belonging to the same young generation have remarkably different distributions. In the field of SN1987A the massive stars are mainly concentrated in a cluster with a radius of 5 pc (the Supernova progenitor was a member of this cluster). The young low mass stars, on the other hand, are more uniformly distributed over the entire field. A similar behavior is also seen in the parallel field. Here, however, there is no clear indication of a clump of massive stars. They are arranged in a linear structure pointing in the direction of the nearby association NGC 2050. In both regions we have observed, the LMC field population is uniformly distributed, as expected, and the average density of these stars is indistinguishable between the two fields.

Star formation history: the Star Formation Rate is found to have monotonically increased, by a factor of at least 6, in the last 8 Gyrs. The star formation activity is still very high at present. In particular, the region around SN1987A has experienced a prominent episode of star formation roughly 12 Myrs ago, i.e. coeval with the birth of the progenitor of SN1987A. Also, the youngest stars in the field, as young as 2 Myrs, appear to be T Tauri stars, characterized by strong Hα emission (EW > 8 Å).

Initial Mass Function: As far as the spatial homogeneity is concerned, we have already mentioned that we do find strong evidence that massive stars are more clustered than low mass ones. Nonetheless, if we average over a suitably large region (roughly 2000 pc^{2} in the case of the field around SN1987A) we find that the concept of IMF between 1 and 7 M_{⊙} is viable and the function is very well represented by a power-law with a logarithmic slope steeper than Γ = −1.55. Unfortunately, we can only place upper bounds on it because it is only possible to place lower limits on the number of young, low mass stars detected through Hα and/or U-band excesses.

Last, but not least, the location of the observed Zero Age Main Sequence is perfectly compatible with a distance to SN1987A of 51.4 ± 1.2 kpc (Panagia, 1998).

PhD thesis completed at the Scuola Normale Superiore di Pisa, Italy on October, 5^{th} 1998 under the direction of Professor Giuseppe Bertin and Professor Nino Panagia. For copies, contact: mromanie@eso.org

Job Opportunity

Faculty Position at the University of Concepción, Chile

The astronomy group of the Physics Department of the Universidad de Concepción, Chile is soliciting applications for a Full Professor Position in observational astronomy. he successful candidate will have full access to all telescopes operating in Chile, including VLT, Magellan 1 and Gemini South, under the 10 percent observing time guaranteed to Chilean astronomers; there is furthermore privileged access to the Gemini Hawaii telescope. The Universidad de Concepción is the most important Chilean university outside Santiago and has recently implemented a strong research program in astronomy and astrophysics. Details about members of our astronomy group, and their research interests, can be found at http://coma.cfm.udec.cl/.

Applicants must be able to assume the position by September 1999. However, the position can be filled as early as April 1999. The position is for three years. Salary is US $ 60,000 per year. After finishing the three year term, there is a possibility of a permanent position at the Universidad de Concepción with a new contract. We invite applications from researchers of all fields of observational astronomy and astrophysics; however, priority will be given to individuals whose research interests have some overlap with the research currently conducted at our Department. Besides a strong research activity, the successful candidate will assume some teaching responsibilities, at a reasonable level. Knowledge of Spanish is advantageous, but not indispensable. Applicants should send a curriculum vitae, bibliography, a statement of research interests, and arrange for three reference letters to be sent to:

Prof. Wolfgang Gieren

Astronomy Group

Departamento de Física

Universidad de Concepción

Casilla 4009

Concepción, Chile

Applicant review will begin March 15, 1999, and continue until the position is filled.

Postdoctoral Position at the University of Concepción, Chile

The astronomy group of the Physics Department of the Universidad de Concepción, Chile is soliciting applications for a Postdoctoral Position. The successful applicant will work with Dr. Wolfgang Gieren on the local calibration of the extragalactic distance scale using Cepheid variables in the Magellanic Clouds and other nearby galaxies, and will have access to several data sets, including near-infrared images of Cepheids in Magellanic Cloud clusters. He/she will also have full access to all telescopes operating in Chile, including VLT, Magellan 1 and Gemini South under the 10 percent observing time guaranteed to Chilean astronomers. The Universidad de Concepción is the most important Chilean university outside Santiago and has recently implemented a strong research program in astronomy and astrophysics. Details about the infrastructure available to the astronomy group can be found at http://coma.cfm.udec.cl/.

Applicants must be able to assume the position by September 1999. However, the position can be filled starting as early as March 1999. The position is for two years, with the possibility of a merit-based extension for a third year. Salary is US $ 30,000 per year. Applicants should have a sound experience in the reduction and analysis of CCD data. Experience with PSF photometry would be advantageous, as well as experience with the reduction of Echelle spectra. Applicants should send a curriculum vitae, bibliography, a statement of research interests, and arrange for three reference letters to be sent to:

Prof. Wolfgang Gieren

Astronomy Group

Departamento de Física

Universidad de Concepción

Casilla 4009

Concepción, Chile

Applicant review will begin March 15, 1999, and continue until the position is filled.

Meeting Announcements

AAS Topical Session:

Shells in the Global ISM

This topical session will be held at the Centennial Meeting of the AAS, Chicago, IL, 30 May – 3 June, 1999. The session will focus on understanding the role of shells in the global ISM of galaxies. Shells and superbubble activity often dominate the kinematics and energy balance in the ISM, which in turn strongly influences star formation activity and galaxy evolution. How do shells mediate and/or regulate the interaction of hot and cool gas, and the dispersion of heavy elements? What mechanisms create shells and structure in the ISM? What are the relative roles of the different types of shells in the multi-phase ISM of different galactic types? How and when does blowout occur, and does material escape from galaxies?

The oral component of this session will take place during the morning and early afternoon on Wednesday, June 2. There will be three themes addressed by the talks:

• Overview of the Multiphase ISM and Structure vs. Hubble Type

• Structure Formation Mechanisms

• Global Effect of Structure vs. Hubble Type

The preliminary list of speakers is: J.N. Bregman, B.G. Elmegreen, A. Ferrara, T.M. Heckman, C. Heiles, J.S. Gallagher, M.-M. Mac Low, R.J. Rand, R.L. Shelton, J.M. Shull, G. Tenorio-Tagle, E. Vazquez-Semadeni, K. Wada, R.A.M. Walterbos, and Q.D. Wang.

Time constraints restrict the oral component to only invited presentations. However, we welcome contributions for the affiliated poster session, to be displayed on the same day. To join the email list for this session, and for instructions on submitting posters, please contact one of us.

Sally Oey (oey@stsci.edu)

You-Hua Chu (chu@astro.uiuc.edu)

Elias Brinks (ebrinks@coyotl.astro.ugto.mx)

AAS Topical Session:

Extragalactic Star Clusters: From Ancient Globulars to Young Super Star Clusters

Time and Venue

This one-day topical session with invited review and highlight talks, a final panel discussion, and a concurrent poster session will be held on June 1 during the AAS Centennial Meeting in Chicago.

Scientific background

Star clusters are observed in a wide range of environments ranging from the high-mass end of the galaxy luminosity function (cDs, ellipticals) to its lowest-mass members (dwarf irregulars, dwarf spheroidals). As luminous single-burst populations, star clusters are unique and long-lasting chronometers of major star formation episodes even at large distances. The preferred formation epoch of globular clusters appears to have been at the earliest times of galaxy formation, while open clusters and OB associations have formed (and dispersed) in gas-rich galaxies at all times. Details of the preconditions for globular cluster formation are not yet understood. Globular cluster formation at later times seems to be linked to galaxy interactions and mergers.

Topics

Our topical session aims at exploring the diversity and common properties of star clusters ranging from ancient globulars to young starburst clusters in different environments. We want to bring together researchers working on the many different aspects of extragalactic star clusters in order to gain a comprehensive picture of star cluster formation and evolution and their role in the evolutionary framework of their parent galaxies. The topical session consists of invited talks that focus on the observational and theoretical aspects of the following topics:

1. Properties of resolved extragalactic star clusters in the Local Group in the context of its evolutionary history 2. Young star cluster populations, starburst clusters, and the formation of globular clusters

3. Observational properties of globular cluster systems

4. Formation scenarios for globular cluster systems

Posters

We invite contributions for a concurrent poster session. Posters will be summarized in the final session, and five of the presentors will be asked to give brief presentations of their posters.

Program and speakers

Our speakers are: Côté (CalTech), Drissen (Laval), Elmegreen (IBM), Geisler (Concepción), Grebel (Washington), Ho (OCIW), Hodge (Washington), Johnson (UCSC), Kennicutt (U Arizona/Steward), Kissler-Patig (ESO), McLaughlin (Berkeley), Miller (Leiden), Sarajedini (Wesleyan), Schweizer (DTM), Zepf (Yale), Zinn (Yale).

The schedule and the titles of the talks are posted at the following URL:

              http://www.astro.washington.edu/grebel/topical.html

Contacts/Organizing Committee

Eva K. Grebel  University of Washington  grebel@astro.washington.edu
Doug Geisler  Universidad de Concepción  doug@noao.edu
Ata Sarajedini  Wesleyan University  ata@astro.wesleyan.edu
Markus Kissler-Patig European Southern Observatory mkissler@eso.org

We hope to see you in Chicago!